OpenStudy (anonymous):

solve the seperable equation: dy/dx= (sec^2)y/1+x^2 im not sure how to reach the final answer

OpenStudy (anonymous):

cos^2(y)dy=(1+x^2)dx 1/2+cos(2y)/4=x+x^3/3 +c

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

it says the answer is 2y + sin(2y) = 4 arctan x +C

OpenStudy (anonymous):

dy/dx= (sec^2)y/1+x^2 dy/sec^2)y= (1+x^2)dx (cos y)^2 dy=(1+x^2)dx integrating both,sides 2y + sin(2y) = 4 arctan x +C

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sorry I read question wrong cos^2(y)dy=(1/1+x^2)dx y/2+sin(2y)/4=aractan(x)+C 2y+sin(2y)=arctan(x)+C